Calculate the momentum. of a body which is moving with a velocity of 36 kaffer and having mass 29 kg. 8 kg..


1
Expert's answer
2021-07-05T08:43:08-0400
"p=mv\\p=29(36)=1044\frac{kg\cdot m}{s}"